OpenStudy (anonymous):

what is the simplest form of the expression sqrt2-sqrt6 over sqrt2+sqrt6

OpenStudy (anonymous):

\[\sqrt{2}-\sqrt{6} \div \sqrt{2} + \sqrt{6}\]

OpenStudy (anonymous):

i think its 1

OpenStudy (kc_kennylau):

you are kidding me @chinenyeogueri

OpenStudy (kc_kennylau):

multiply the conjugate of the denominator to the denominator and the numerator :)

OpenStudy (anonymous):

@kc_kennylau damn i said i think get the flutter out

OpenStudy (kc_kennylau):

?

OpenStudy (kc_kennylau):

\[\frac{\sqrt2-\sqrt6}{\sqrt2+\sqrt6}=\frac{(\sqrt2-\sqrt6)(\sqrt2-\sqrt6)}{(\sqrt2+\sqrt6)(\sqrt2-\sqrt6)}=\frac{(\sqrt2-\sqrt6)^2}{2-6}\]